Jeanne Lewis is attempting to evaluate two possible portfolios consisting of the same five assets but held in different proportions. She is particularly interested in using beta to compare the risk of the portfolios​ and, in this​ regard, has gathered the following​ data

Portfolio Weights  
Asset   Asset Beta   Portfolio A   Portfolio B
1 1.33   12%   31%
2 0.71   33%   11%
3 1.23   10%   25%
4   1.15   5%   17%
5 0.86   40%   16%
Total   100%   100%

a. Calculate the betas for portfolios A and B.

b. Portfolio (blank) is slightly less risky than the market, while portfolio (blank) is more risky than the market. Portfolio's (blank) return will move more than portfolio (blank's) for a given increase or decrease in market risk. Portfolio (blank) is the more risky portfolio.
Answer #1

Answer (a):

Beta of Portfolio A = 1.33 * 12% + 0.71 * 33% + 1.23 * 10% + 1.15 * 5% + 0.86 * 40% = 0.92

Beta of Portfolio B = 1.33 * 31% + 0.71 * 11% + 1.23 * 25% + 1.15 * 17% + 0.86 * 16% = 1.13

Hence:

Beta of Portfolio A = 0.92

Beta of Portfolio B = 1.13

Answer (b):

Portfolio A is slightly less risky than the market, while portfolio B is more risky than the market. Portfolio B's return will move more than portfolio A's for a given increase or decrease in market risk. Portfolio B is the more risky portfolio.